# Locret: Enhancing Eviction in Long-Context LLM Inference with Trained Retaining Heads

### v1

@@ 3.2 Locret  Framework
Locret is a training-based KV cache compression framework that works in conjunction with chunked prefill. As illustrated in Figure 1, Locret operates in two stages: training and inference. In the training stage, we modify the original LLM by appending a retaining head $\mathbf{R}$ to each attention module. We then train the retaining heads $\mathbf{R}$ while keeping the LLM backbone frozen. During chunked prefill inference, the retaining heads $\mathbf{R}$ are used to calculate the importance of each cache unit in the chunk. We retain the cache units with higher scores, along with stabilizers (i.e., the last tokens), in the cache pool located in GPU memory. Through this process, the retaining heads $\mathbf{R}$ learn and predict the patterns discovered by existing methods, as detailed in Appendix G.
From a mathematical perspective, cache eviction is performed by assigning each cache unit an importance score that reflects its influence on the calculation of subsequent tokens. We refer to this estimation as the causal importance score (CIS) since it is computed in a causal manner. The CIS of cache unit $k$ is calculated as $s_{k}=S(c_{1},c_{2},\cdots,c_{k})$. By applying top-$b$ sparse attention based on the CIS, we can ensure that the trace (i.e. retaining and eviction) of each cache unit can be fit within a cache with a given memory budget. Further details can be found in Appendix F.
However, since not all tokens can be stored in the cache simultaneously, calculating the actual CIS on-chip is impractical. Instead, we use a heuristic approximation for CIS, defined as follows: $\tilde{s}_{k}=S(\tilde{c}_{p_{1}},\tilde{c}_{p_{2}},\cdots,\tilde{c}_{p_{b^{\prime}}})$, where $b$ is the cache budget, all $\tilde{c}_{p_{i}}$ are approximated cached units, and $b^{\prime}\leq b$. We hypothesize that if the scoring function for causal importance is sufficiently accurate, it will consistently select the most critical cache units, resulting in a negligible difference between the heuristic and the actual score. Thus, we use the terms heuristic CIS and actual CIS interchangeably.
3.3 Training the Retaining Heads
@@
In this section, we introduce Locret’s model architecture modifications and the corresponding training recipe. We add additional parameters to compute the CIS $s_{k}$ (or $\tilde{s}_{k}$ for on-chip inference) with respect to all previous cache units. Specifically, we inject a retaining head, consisting of a small MLP, into each layer. The retaining head predicts the CIS for each head of the corresponding layer based on the concatenation of $[\mathbf{Q},\mathbf{K},\mathbf{V}]$. Formally, with a slight abuse of notation, let the retaining head for layer $i$ be denoted as $\mathbf{R}$. The CIS at head $j$ of layer $i$ is then calculated as: $\mathbf{\tilde{S}}=\mathbf{R}([\mathbf{Q},\mathbf{K},\mathbf{V}])=\sigma([\mathbf{Q},\mathbf{K},\mathbf{V}]\mathbf{W}_{1})\mathbf{W}_{2}$. In this equation, $\mathbf{W}_{1}\in\mathbb{R}^{(d_{m}+2d_{kv})\times d_{\mathbf{R}}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{d_{\mathbf{R}}\times\frac{h}{g}}$ are the tunable parameters of $\mathbf{R}$, $\sigma$ is the activation function and $\mathbf{\tilde{S}}[k]_{j}$ is the predicted CIS of the $k$-th token at head $j$ of layer $i$. This architecture implies that the importance estimation for a single head is not performed in isolation but rather considers all heads together. Note that for GQA models, there are only $h/g$ output values corresponding to the number of heads in the KV cache.
@@
We train the retaining head $\mathbf{R}$s on a small Question-Answer (QA) supervised fine-tuning (SFT) dataset, where each entry consists of a single prompt and one answer. We define the CIS $s_{k}$ for the $k$-th token as the maximum attention score, before softmax, from all the answer tokens toward the $k$-th token. Formally, for the $k$-th token at head $j$ of layer $i$, we approximate the predicted value $\mathbf{\tilde{S}}[k]_{j}^{(i)}$ to the ground truth $\mathbf{S}[k]_{j}^{(i)}:=\max_{p}\left(\mathbf{Q}_{j}^{(i)}\mathbf{K}_{j}^{(i)T}\right)_{p,k}$, where $n_{q}(d)\leq p\leq n_{q}(d)+n_{a}(d)$, and $n_{q}(d)$ and $n_{a}(d)$ represent the lengths of the prompt and answer in data $d$, respectively. For an MHA model with $L$ layers and $h$ heads, the training objective is described in Equation 2. For GQA models, we take the maximum attention score before softmax across different query heads within the same group as the ground truth for the corresponding KV head.
@@
$\operatorname*{arg\,min}_{\mathbf{W}_{1}^{(i)},\mathbf{W}_{2}^{(i)},i=1,2\cdots,L}\mathbb{E}_{d\in\mathcal{D}}\left[\sum_{i=1}^{L}\sum_{j=1}^{h}\sum_{k=1}^{n_{q}(d)}\mathcal{L}\left(\mathbf{\tilde{S}}[k]_{j}^{(i)},\mathbf{S}[k]_{j}^{(i)}\right)\right]$ (2)
@@
$\mathcal{L}\left(\mathbf{\tilde{S}}[k]_{j}^{(i)},\mathbf{S}[k]_{j}^{(i)}\right)=\text{Smooth-}\mathcal{L}_{1}\left(\mathbf{\tilde{S}}[k]_{j}^{(i)},\mathbf{S}[k]_{j}^{(i)}\right)+\alpha\mathcal{L}_{2}\left(\mathbf{\tilde{S}}[k]_{j}^{(i)},\mathbf{\tilde{S}}[k+1]_{j}^{(i)}\right)$ (3)
Table 1: L-Eval scores of Locret trained on various datasets.
Dataset LongAlpaca LongAlign Anti-Haystack
L-Eval 21.33 22.00 20.72
From our observations, the training of Locret exhibits strong robustness. The performance variations shown in Figure 2 and Table 2 are minimal, despite changes in $d_{\mathbf{R}}$ and the dataset. Details can be found in Appendix D. Training statistics, including loss dynamics, are recorded in Appendix H.
@@ 3.4 Inference Implementation of Locret
Output: Generated tokens $x_{gen}$
// Leave the last $n_{loc}$ out to make sure they are not evicted.
chunk_positions $\leftarrow$ split_chunk(00, $x$.length() $-n_{loc}$, $B$)
K_cache, V_cache, score_cache $\leftarrow$ [], [], []
for chunk $\in$ chunk_positions do
      begin_pos, end_pos $\leftarrow$ chunk.begin_pos, chunk.end_pos
[…]
      K_cache $\leftarrow$ Concat(K_cache, K_chunk)
      V_cache $\leftarrow$ Concat(V_cache, V_chunk)
      score_cache $\leftarrow$ Concat(score_cache, score_chunk)
      if chunk is not the last chunk then
           // Keep the last $n_{s}$ caches to maintain higher context continuity.
           score_cache[score_cache.length()-$n_{s}$:score_cache.length()] $\leftarrow$+$\infty$ + ∞
           end if
          indices $\leftarrow$ top-$b$(score_cache).indices
           K_cache, V_cache, score_cache = K_cache[indices], V_cache[indices], score_cache[indices]
           end for
          K_cache, V_cache, score_cache $\leftarrow\mathbf{M}$($x$[$x$.length()$-n_{loc}$:$x$.length()], K_cache, V_cache)
           $x_{gen}\leftarrow\mathbf{M}$.generate(K_cache, V_cache)
           return $x_{gen}$
Algorithm 1 Locret Inference
During the inference stage, we use the chunked prefill pattern and perform cache eviction based on the predicted CIS. Since the predicted value $\mathbf{\tilde{S}}[k]_{j}^{(i)}$ depends solely on $\mathbf{Q}[k]_{j}^{(i)}$, $\mathbf{K}[k]_{j}^{(i)}$, and $\mathbf{V}[k]_{j}^{(i)}$, and because attention in decoder-only models is causal, $\mathbf{\tilde{S}}[k]_{j}^{(i)}$ remains consistent once calculated. Thus, we store the KV cache units along with their corresponding causal importance values. When the cache is full, we evict the units with lower causal importance values, as they are deemed less useful for future computations. Cache eviction introduces context discontinuity, meaning some cache units at certain positions may be absent. This can degrade generation quality and increase the error between the predicted and accurate CIS, as LLMs are typically not trained on such contexts. To mitigate this, we retain the last tokens of the current chunk at each step of the chunked prefill process, ensuring a local and continuous context to minimize errors. To demonstrate the effectiveness of this design, we perform an ablation study on the length of stabilizers $n_{s}$, shown in Figure 3. Smaller $n_{s}$ results in severe performance degradation, and the model fails entirely when stabilizers are absent, as context discontinuity leads to instability in CIS prediction, causing errors in cache eviction and amplifying errors in hidden states. More details are discussed in Appendix E.
We maintain a cache pool with a capacity of $b$ cache units, discarding units that exceed this limit. For each chunk, the model processes the chunked input tokens alongside the current cache pool. The newly generated KV pairs and their predicted scores are then concatenated with the existing cache. Once the cache pool is full, only the $b$ cache units with the highest CIS values are retained. At each chunked prefill step, except for the final step, we retain the stabilizers, i.e. the last $n_{s}$ cache units. Additionally, we do not compress the last $n_{loc}$ tokens of the prompt, as they are critical for maintaining high generation quality due to their strong correlation with the query. Finally, the answer is generated according to the compressed KV cache. Algorithm 1 provides the pseudocode for Locret, where we formally describe the Locret inference process.
The GPU memory usage during Locret inference can be effectively bounded. GPU memory for KV cache storage is limited to $\mathcal{O}(b+n_{loc})$, and the runtime memory usage of the attention mechanism is bounded by $\mathcal{O}(B\times(b+B+n_{loc}))$. For comparison, while processing an input with $n$ tokens, full attention prefill requires $\mathcal{O}(n)$ for KV cache storage and $\mathcal{O}(n^{2})$ for runtime memory, whereas chunked prefill requires $\mathcal{O}(n)$ for KV cache storage and $\mathcal{O}(nB)$ for runtime memory consumption.
@@ 4 Experiments
In this section, we present the experiments conducted to evaluate the proposed framework, Locret, aiming to address the following questions:
(Q1) Can Locret obtain better end-to-end task performance compared to popular and competitive long-conetext inference approaches within similar or less peak memory?
(Q2) Can Locret improve inference speed compared to other approaches?
(Q3) Can Locret be combined with other efficient inference methods?
4.1 Experimental Setup
Models and training dataset. We evaluate Locret on two long-context LLMs: Phi-3-mini-128K (Abdin et al., 2024) and Llama-3.1-8B-instruct (Dubey et al., 2024). Both models can process up to 128K context tokens and follow MHA and GQA architectures, respectively. Phi-3-mini-128K has 3.8B parameters, while Llama-3.1-8B-instruct contains 8B parameters, making both suitable for deployment on consumer-grade devices. We inject retaining heads $\mathbf{R}$ into each layer, setting the intermediate size $d_{\mathbf{R}}$ to 1024 for both models. The retaining heads are trained on the LongAlpaca dataset (Chen et al., 2024) for 3000 steps. Training Locret is lightweight, with the tunable parameters comprising 8% and 2.5% of the total for the two models, respectively. The complete training process takes 0.47 and 0.80 GPU hours on a single A800 GPU for each corresponding model. The cache budget $b$ is set to 6000 for Phi-3-mini-128K and 16384 for Llama-3.1-8B-instruct. More details on hyperparameters for both training and inference, as well as the system environment, can be found in Appendix A.
[…]
Baselines. As discussed in Section 2, existing algorithms for memory-efficient long-context inference can be categorized into offloading-based, quantization-based, token-dropping, and sparsity-based methods. For each category, we select one representative method as the baseline. We compare Locret against full attention inference (denoted as FullAttn), InfLLM (Xiao et al., 2024a), KV cache quantization (Turganbay, 2024), SirLLM (Yao et al., 2024), and MInference (Jiang et al., 2024a). For quantization, we use Hugging Face Quanto (Hugging-Face, ) implementation, referring to the 2-bit quantization method as HF-2bits. We omit HF-4bits and benchmark this combination in Section 4.4. We do not include attention pool-based token-dropping methods in this benchmark, as they are orthogonal to our approach; further discussion is provided in Section 4.4. Detailed introductions to the selected baselines can be found in Appendix A. We also discuss the comparison between the trained Locret and the randomly initialized retaining heads $\mathbf{R}$ in Appendix B.
4.2 End-to-end Benchmark
[…]
In the L-Eval tests, all methods show some degree of performance degradation. Nevertheless, Locret achieves the best overall result, obtaining the highest score on most tasks. L-Eval is a shorter but more complex dataset, where SirLLM performs particularly well. Quantization fails on most tasks, resulting in the lowest overall score. Both InfLLM and MInference suffer significant performance drops compared to FullAttn inference. Locret consistently surpasses all competitors.
[…]
Figure 4: Memory Statistics vs. Task Performance. The red lines correspond to the theoretical size of the model weights, while the blue lines represent the total theoretical size of the model weights and the full KV cache without any compression. The purple lines indicate the accuracies of FullAttn. “Total Memory” represents the total memory usage of both GPU and CPU memory.
@@ 4.2 End-to-end Benchmark
4.3 Speed Test on Real Consumer-Grade Devices
@@
R.PassKey is a task where the model retrieves a 5-digit number from a large amount of irrelevant text, a task we believe to be relatively simple for humans. Thus, we consider the task to have failed if the accuracy falls below 95%. As shown in Table 3, aside from the settings that fail on this task, Locret achieves the highest inference speed among all methods that can correctly process R.PassKey. Due to its MHA architecture, Phi-3-mini-128K has a larger KV cache, which leads to failures for both HF-2bits and MInference. Storing the full KV cache on a single 4090 GPU is infeasible, as it requires 48GB of memory. Although the quantized KV cache is reduced to 6GB, the converting processes between representations requires significant GPU memory for intermediate states, resulting in the failure of HF-2bits. While InfLLM can run in memory-limited scenarios, its offloading process slows down inference, with I/O becoming the bottleneck in attention calculation. SirLLM fails due to its inaccurate eviction policy, which cannot correctly identify the 5-digit number. In the GQA model (Llama-3.1-8B-instruct), which has a smaller KV cache, the quantized cache can fit within the GPU memory. However, the quantization and dequantization processes become the bottleneck, leading to significantly slower speeds. The performance of InfLLM, SirLLM, and MInference is similar to that seen with Phi-3-mini-128K. Although MInference benefits from faster encoding speeds, it fails on this task because it cannot process the entire input sequence at once. Locret strikes a balance between inference speed and performance, making it a far more suitable solution for long-context scenarios on consumer-grade devices.
Table 3: Executing R.PassKey on an Nvidia 4090. “tok/s” represents the inference speed, “C.Len” stands for the context length after truncation, and “Acc.” represents task accuracy. “Success” shows whether a setting has $\geq$95accuracy. The highest score among 128K context is marked in bold.

### v2

@@ 3.2 Framework of Locret
Locret is a training-based KV cache eviction framework that works in conjunction with chunked prefill. As illustrated in Figure 2, Locret operates in two stages: training and inference. In the training stage, we modify the original LLM by appending a retaining head $\mathbf{R}$ to each attention module. We then train the retaining heads $\mathbf{R}$ while keeping the LLM backbone frozen. During the chunked prefill inference stage, the retaining heads $\mathbf{R}$ can obtain the importance of each cache unit. We retain the cache units with higher scores, along with stabilizers (i.e., the last tokens), in the cache pool located in GPU memory. Through this process, the retaining heads $\mathbf{R}$ learn and predict the patterns discovered by existing methods, e.g. attention sink in Xiao et al. (2024c) and vertical lines in MInference (Jiang et al., 2024a), as detailed in Appendix O.
The eviction policy assigns each cache unit an importance score reflecting its influence on comprehending subsequent context. This estimation is causal, termed the causal importance score (CIS). The CIS of the $k$-th unit depends only on the preceding units and the $k$-th unit itself. Due to memory constraints, calculating the exact CIS on-chip is impractical. Intuitively, discarding certain KV cache during the attention calculation introduces errors, manifesting as biases in the CIS. However, as long as the importance estimation is sufficiently accurate, the core information can be preserved, and the loss of information in the attention process can be minimized. Therefore, the error introduced in the CIS can be neglected. Please refer to Appendix N for more details.
3.3 Training Retaining Heads
@@
In this section, we introduce Locret’s model architecture modifications and the corresponding training recipe. We add additional parameters to compute the CIS $\mathbf{S}[k]$ for the $k$-th cache unit. Specifically, we inject a retaining head, consisting of a small MLP, into each layer. From our observation, such small MLPs do not slow down model inference, with details elaborated in Appendix M. The retaining head predicts the CIS for each head of the corresponding layer based on the concatenation of $[\mathbf{Q},\mathbf{K},\mathbf{V}]$. Formally, with a slight abuse of notation, let the retaining head for layer $i$ be denoted as $\mathbf{R}$. The CIS at head $j$ is then calculated as: $\mathbf{\tilde{S}}=\mathbf{R}([\mathbf{Q},\mathbf{K},\mathbf{V}])=\sigma([\mathbf{Q},\mathbf{K},\mathbf{V}]\mathbf{W}_{1})\mathbf{W}_{2}$ In this equation, $\mathbf{W}_{1}\in\mathbb{R}^{(d_{m}+2d_{kv})\times d_{\mathbf{R}}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{d_{\mathbf{R}}\times\frac{h}{g}}$ are the tunable parameters of $\mathbf{R}$, and $\sigma$ is the activation function, $\mathbf{\tilde{S}}[k]=\left[\mathbf{\tilde{S}}[k]_{1},\cdots,\mathbf{\tilde{S}}[k]_{h/g}\right]\in\mathbb{R}^{\frac{h}{g}}$ , where $\mathbf{\tilde{S}}[k]_{j}$ is the predicted CIS of the $k$-th token. This architecture implies that the importance estimation for a single head is not performed in isolation but considers all heads.
@@
We train the retaining head $\mathbf{R}$s on a small Question-Answer (QA) supervised fine-tuning (SFT) dataset, where each entry consists of a single prompt and one answer. We define the CIS $\mathbf{S}[k]_{j}$ for the $k$-th token at head $j$ as the maximum attention score, before softmax, from all the answer tokens toward the $k$-th token. Formally, given a training instance $d$, for the $k$-th token at head $j$ of layer $i$, we approximate the predicted value $\mathbf{\tilde{S}}[k]_{j}^{(i)}$ to the ground truth $\mathbf{S}[k]_{j}^{(i)}:=\max_{p}\left(\mathbf{Q}_{j}^{(i)}\mathbf{K}_{j}^{(i)T}\right)_{p,k}$ , where $n_{q}(d)<p\leq n_{q}(d)+n_{a}(d)$, and $n_{q}(d)$ and $n_{a}(d)$ represent the lengths of the prompt and answer in data $d$, respectively. For an MHA model with $L$ layers and $h$ heads, the training objective is described in Equation 2. For GQA models, we take the maximum attention score before softmax across different query heads within the same group as the ground truth for the corresponding KV head.
@@
$\underset{\mathbf{W}_{1}^{(i)},\mathbf{W}_{2}^{(i)},i=1,2\cdots,L}{\text{argmin}}\mathbb{E}_{d\in\mathcal{D}}\left[\sum_{i=1}^{L}\sum_{j=1}^{h}\sum_{k=1}^{n_{q}(d)}\mathcal{L}\left(\mathbf{\tilde{S}}[k]_{j}^{(i)},\mathbf{S}[k]_{j}^{(i)}\right)\right]$ (2)
@@
[…]
$+\alpha\mathcal{L}_{2}\left(\mathbf{\tilde{S}}[k]_{j}^{(i)},\mathbf{\tilde{S}}[k+1]_{j}^{(i)}\right)$
From our observations, the training of Locret exhibits strong robustness. Despite changes in $d_{\mathbf{R}}$ and the dataset, the performance variations shown in Figure 6 and Table 6 are minimal. Details can be found in Appendix I. Training statistics and loss dynamics are recorded in Appendix P.
@@ 3.4 Inference Implementation of Locret
During the inference stage, we use the chunked prefill pattern and perform cache eviction based on the predicted CIS. Since the predicted CIS does not rely on subsequent tokens, it remains consistent once calculated. Thus, we store the KV cache units along with their corresponding causal importance values. When the cache is full, we evict the units with lower causal importance values, as they are deemed less useful for future computations. Such eviction is performed during chunked prefill. When processing a new chunk, we first compute its KV cache, concatenate it with the previously retained cache, and evict redundant units to adhere to the budget size. Note that we cache the pre-RoPE KV cache and reassign continuous position embeddings from the beginning to enhance context continuity.
Cache eviction introduces context discontinuity, meaning some cache units at certain positions may be absent, which can degrade generation quality. To mitigate this, we retain the last $n_{s}$ tokens of the current chunk, named as the stabilizers, at each step of chunked prefill, ensuring a local and continuous context to minimize errors. As shown in Figure 3, smaller $n_{s}$ results in severe performance degradation, and the model fails entirely when stabilizers are absent, as context discontinuity leads to instability in CIS prediction, causing errors in cache eviction and amplifying errors in hidden states. More details are discussed in Appendix L. We provide a pseudocode of Locret inference in Algorithm 1.
@@ 4 Experiments
We conduct experiments to evaluate whether Locret can address the following questions:  (Q1) Can Locret obtain better end-to-end task performance compared to popular and competitive long-context inference methods using similar or less peak memory?  (Q2) Is Locret able to achieve a faster inference speed on consumer-grade devices?  (Q3) How can Locret process query-driven tasks?
@@
Figure 4: Memory Statistics vs. Task Performance. The red lines correspond to the theoretical size of the model weights, while the blue lines represent the total size of the model weights and the full KV cache without any compression. The purple lines indicate the accuracies of FullAttn. “Total Memory” represents the total memory usage of both GPU and CPU.
4.1 Experimental Setup
Model and Dataset Settings. We conduct experiments on two long-context LLMs: Phi-3-mini-128K (Abdin et al., 2024) and Llama-3.1-8B-instruct (Dubey et al., 2024). Both models can process up to 128K context tokens and follow MHA and GQA architectures, respectively. The parameter sizes of these two models are also suitable for deployment on consumer-grade devices. We inject retaining heads $\mathbf{R}$ into each layer of these two models, and the intermediate size $d_{\mathbf{R}}$ is 1,024. The retaining heads are trained on the LongAlpaca dataset (Chen et al., 2024) for 3,000 steps , with a 5e-4 learning rate, 10,240 sequence length, and $\alpha$ set to 2.5e-3. Training Locret is lightweight, with the tunable parameters comprising 8% and 2.5% of the total for the two models, respectively. The complete training process takes 0.47 and 0.80 GPU hours on an A800 GPU for each corresponding model. More important hyperparameters are listed in Table 4. More details on hyperparameters and system environments can be found in Appendix A.
@@ 4.1 Experimental Setup
Baselines. As discussed in Section 2, existing algorithms for memory-efficient long-context inference can be categorized into offloading-based, sparsity-based, quantization-based, and token-dropping methods. For each category, we select one representative method as the baseline. We compare Locret against full attention inference (denoted as FullAttn), InfLLM (Xiao et al., 2024a), MInference (Jiang et al., 2024a), KV cache quantization (Turganbay, 2024), and SirLLM (Yao et al., 2024). For quantization, we use HuggingFace Quanto (Hugging-Face, ) implementation, referring to the 2-bit quantization method as HF-2bits. We omit HF-4bits and benchmark the combination with Locret in Section H. We do not include attention pooling-based token-dropping methods in this benchmark, as they are orthogonal to our approach, and further discussion about this is provided in Section H. Detailed introductions to the selected baselines can be found in Appendix A. We also discuss the comparison between the trained Locret and the randomly initialized retaining heads $\mathbf{R}$ in Appendix E.
4.2 End-to-end Benchmark
[…]
(2) On L-Eval, all methods show performance degradation. Nevertheless, Locret achieves the best overall performance, obtaining the highest scores on most tasks. L-Eval is a shorter but more complex dataset, where SirLLM performs particularly well. Quantization fails on most tasks. Both InfLLM and MInference suffer significant performance drops compared to FullAttn inference. Locret consistently surpasses all competitors.
We report memory consumption in Figure 4, showing:
[…]
(2) For Phi-3-mini-128K, which has a larger KV cache, InfLLM and MInference exhibit higher memory consumption since they store the full KV cache. Although other methods have similar memory footprints, Locret uses the least memory and achieves the best overall performance.
(3) For Llama-3.1-8B-instruct, whose full KV cache is smaller, the memory bottleneck shifts to the runtime computational memory for attention and other calculations. All methods exhibit similar memory footprints, with Locret delivering the best overall performance.
In summary, our experiments demonstrate that Locret is both effective and efficient, outperforming all baselines on multiple datasets and models while using less GPU memory.
Table 2: Executing R.PassKey on an NVIDIA 4090. “tok/s” represents the inference speed, “C.Len” stands for the context length after truncation, and “Acc.” represents task accuracy. The highest score among 128K context is marked in bold.